OpenStudy (anonymous):

Find the slope of the following graph.... Find the slope of the following graph.... @Mathematics

OpenStudy (anonymous):

Here is the attatchment.

OpenStudy (saifoo.khan):

3

OpenStudy (anonymous):

Why?

OpenStudy (saifoo.khan):

use the formula. :p i did visually.

OpenStudy (anonymous):

deltay/deltax

OpenStudy (anonymous):

choose 2 points (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

Here is another.

OpenStudy (anonymous):

I put 0 because that's what I thought the answer was.

OpenStudy (saifoo.khan):

Undefined. :P

OpenStudy (anonymous):

Well isn't that the same thing?

OpenStudy (saifoo.khan):

No.

OpenStudy (saifoo.khan):

0/1 = 0 1/0 = undefined.

OpenStudy (anonymous):

Ok Can you help me with another?

OpenStudy (saifoo.khan):

No,

OpenStudy (saifoo.khan):

lol, jk. post it

OpenStudy (anonymous):

Hold on I gotta make a screenshot of the program.

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

Ok here it is.

OpenStudy (saifoo.khan):

\[5 x+3 y-34 = 0\]

OpenStudy (anonymous):

That doesn't work. Lol if you looked at the picture that many numbers does not fit.

OpenStudy (saifoo.khan):

5x + 3y = 34

OpenStudy (anonymous):

Thanks a lot. I needed that! Lol

OpenStudy (anonymous):

Are you still in High School?

OpenStudy (saifoo.khan):

No, i just passed 11th. i help online college/school people.. the study in US is very easy!

OpenStudy (anonymous):

So you live out of Country?

OpenStudy (saifoo.khan):

No, i live in another country. hehe. are u in HS?

OpenStudy (anonymous):

Yeah... I am a Junior.

OpenStudy (anonymous):

Hold on I gotta get something...

OpenStudy (saifoo.khan):

Ohh.. kk

OpenStudy (anonymous):

Haha I am back... Yeah I am in High School. I was adopted from Romania when I was eight. Lol

OpenStudy (saifoo.khan):

wow, cool. =D im from Pak.

OpenStudy (anonymous):

Cool, small World? Huh. I have always had problems with Math... no learning languages is easy for me. Lol this is my fifth language.

OpenStudy (saifoo.khan):

Haha, great.

OpenStudy (anonymous):

Lol. So do you do tutoring?

OpenStudy (saifoo.khan):

Yea, i teach people online.. and do tutoring stuff. HS / Colleges both.

OpenStudy (anonymous):

Ok How do you tutor?

OpenStudy (saifoo.khan):

through skype, sometimes i use whiteboards to explain people steps and stuff.

OpenStudy (anonymous):

How old are you?

OpenStudy (saifoo.khan):

i turned 17 this oct.

OpenStudy (anonymous):

So if I needed help with something, I can do that? Have you looked at Google Voice... they have a free cell number service through the internet.

OpenStudy (saifoo.khan):

Yea, i do but that dosnt work from Pak anymore. we can also use skype, it's free as well.

OpenStudy (anonymous):

Yeah I hate Skype... Lol do you have gmail?

OpenStudy (anonymous):

What is the tradition their for Thanksgiving?

OpenStudy (anonymous):

Or Christmas?

OpenStudy (anonymous):

We can chat through GMail.

OpenStudy (saifoo.khan):

einstein.saifoo@gmail.com we dont really have thanksgiving here. Christmas is celebrated by Christians here in Pak. sometimes we celebrate them with as well as i have friends at school.

OpenStudy (saifoo.khan):

how about fb?

OpenStudy (anonymous):

Yea I don't get on Facebook that often.

OpenStudy (saifoo.khan):

k, gmail works fine then

OpenStudy (anonymous):

Haha, thanks lol. What is your email?

OpenStudy (saifoo.khan):

i wrote that above, einstein.saifoo@gmail.com

OpenStudy (anonymous):

Thanks.

OpenStudy (saifoo.khan):

uw.

OpenStudy (anonymous):

What is UW?

OpenStudy (saifoo.khan):

uw = you welcome.

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

Oh... Hahah

OpenStudy (saifoo.khan):

hehe.

OpenStudy (anonymous):

Multitasking is fun.. lOL

OpenStudy (saifoo.khan):

Haha, that is Multi-Fun. xD

OpenStudy (anonymous):

Hahaha... Lol